OpenStudy (anonymous):

How many grams of NaOH can be formed when 43.0 grams of Na2O react with 27.0 grams of H2O? Na2O + H2O yields NaOH

OpenStudy (abhisar):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (abmon98):

First,construct a balanced equation Na2O + H2O yields NaOH Write down the number of atoms that you have on each side of the equation. On the reactant side: Na:2 O:2 H:2 On the Product Side: Na:1 O:1 H:1 Add a coeffcient of 2 to NaOH Na2O + H2O yields 2NaOH Number of Moles=Mass(g)/MolarMass(g/mol) 43/(23+16)=1.10 moles of Na2O 27/18= 1.5 moles of H2O Mole Ratio of Na2O to H2O are 1:1 Calculate the moles of H2O based on the number of moles of Na2O 1.10 , H2O should contain 1.10 moles but the actual number of moles present is 1.5 and this indicates that H2O is in excess and Na2O is the limiting reagent. The limiting reagent is the reactant that is completely used up in a reaction and thus determines when the reaction stops. so we use Na2O moles to find out the grams produced of NaOH Look at the Mole Ratio of Na2O to NaOH its 1:2 therfor we should multiply 1.10*2=2.20 moles of NaOH are produced Use Number of Moles=Mass(g)/Molar Mass(g/mol) Arrange this equation to find out the mass Mass(g)=Molar Mass(g/mol)*Number of Moles 2.20*(23+16+1)=Mass(g)

OpenStudy (abhisar):

Did you know you can give medals to people who help you with questions? Just click the \(\color{blue}{\text{blue "best response" button}}\) to show appreciation for the people who help you.

OpenStudy (anonymous):

I'm still confused at the end..

OpenStudy (anonymous):

@Abmon98

OpenStudy (anonymous):

How many grams of H2 are needed to react completely with 44.1 grams of C? C + H2 yields CH4

OpenStudy (anonymous):

How far along can you go into the problem on your own? Lets work on it step by step.

OpenStudy (anonymous):

Do you know how to balance the equation? This is always the first step with these kinds of problems.